# Constructing Per-Shot Bitrate Ladders using Visual Information Fidelity

### v1

@@ V-H Constructing Per-Shot Bitrate and Quality Ladders
$q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-15.311/15.217$ $3.042/3.383$ $12.546/14.353$ $-2.422/2.534$ $0.767$ $0.667$ $0.533$
$\text{LLF}_{3}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.546/16.722$ $4.599/4.107$ $0.615/3.546$ $-0.05/0.786$ $1.0$ $0.9$ $0.9$
$\text{VIFF}_{1}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-18.337/17.154$ $4.064/4.225$ $3.274/5.763$ $-0.513/1.069$ $0.833$ $0.767$ $0.6$
$\text{VIFF}_{2}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-17.3/17.353$ $3.728/4.149$ $4.524/5.619$ $-0.747/1.024$ $0.833$ $0.733$ $0.533$
$\text{VIFF}_{3}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-17.905/16.652$ $3.905/4.149$ $3.829/4.074$ $-0.597/0.715$ $0.867$ $0.8$ $0.633$
$\text{VIFF}_{4}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-17.807/17.625$ $3.988/4.293$ $3.676/6.005$ $-0.606/1.108$ $0.833$ $0.733$ $0.6$
$\text{VIFF}_{5}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-18.523/16.967$ $4.091/4.229$ $2.735/4.306$ $-0.42/0.867$ $0.9$ $0.767$ $0.533$
$\text{VIFF}_{6}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-18.677/16.35$ $4.121/4.001$ $3.026/4.971$ $-0.507/1.059$ $0.867$ $0.833$ $0.7$
$\text{VIFF}_{7}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-16.877/17.021$ $3.717/4.231$ $5.288/6.32$ $-0.758/1.086$ $0.8$ $0.7$ $0.567$
$\text{VIFF}_{8}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-16.851/16.555$ $3.695/4.08$ $5.287/6.547$ $-0.78/1.15$ $0.833$ $0.7$ $0.433$
$\text{VIFF}_{9}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-19.272/16.235$ $4.295/4.031$ $2.609/5.824$ $-0.423/1.244$ $0.933$ $0.933$ $0.833$
$\text{LLF}_{3}$, $\text{VIFF}_{1}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.514/16.74$ $4.549/4.054$ $0.471/3.69$ $-0.015/0.822$ $0.933$ $0.933$ $0.867$
$\text{LLF}_{3}$, $\text{VIFF}_{2}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.138/16.605$ $4.437/3.985$ $0.94/3.834$ $-0.103/0.847$ $1.0$ $0.9$ $0.833$
$\text{LLF}_{3}$, $\text{VIFF}_{3}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.5/16.748$ $4.489/4.041$ $0.545/3.815$ $-0.022/0.832$ $0.967$ $0.933$ $0.833$
$\text{LLF}_{3}$, $\text{VIFF}_{4}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.847/16.602$ $\mathbf{4.654}/4.108$ $0.182/3.409$ $0.026/0.766$ $0.967$ $0.967$ $0.9$
$\text{LLF}_{3}$, $\text{VIFF}_{5}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.606/16.645$ $4.498/4.007$ $0.445/3.576$ $0.004/0.79$ $0.933$ $0.933$ $0.867$
$\text{LLF}_{3}$, $\text{VIFF}_{6}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.559/16.655$ $4.568/4.128$ $0.602/3.751$ $-0.032/0.83$ $0.967$ $0.933$ $0.9$
$\text{LLF}_{3}$, $\text{VIFF}_{7}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.5/16.677$ $4.545/4.059$ $0.55/3.798$ $-0.021/0.84$ $1.0$ $0.9$ $0.9$
$\text{LLF}_{3}$, $\text{VIFF}_{8}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $-20.632/16.79$ $4.558/4.043$ $0.405/3.826$ $-0.001/0.836$ $0.967$ $0.9$ $0.833$
$\text{LLF}_{3}$, $\text{VIFF}_{9}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ $\mathbf{-20.951}/16.593$ $4.63/4.076$ $-0.071/3.127$ $0.088/0.729$ $\mathbf{1.0}$ $\mathbf{0.967}$ $\mathbf{0.9}$
$\text{LLF}_{1}$ (Cross-Over VMAFs) $-20.863/16.728$ $4.545/4.247$ $\mathbf{-0.129}/3.304$ $\mathbf{0.102}/0.741$ $0.933$ $0.967$ $0.867$
@@
The performances of the per-shot bitrate ladders and quality ladders, constructed using the various described methods, are presented in Tables IX and X, respectively. These are compared against a fixed bitrate ladder [3] and reference exhaustive bitrate ladders. In Table IX, all rows, except the last one, show the performances of bitrate ladders constructed using quality prediction methods. The last row presents the performance of bitrate ladders constructed by predicting cross-over bitrates. Similarly, in Table X, all rows, barring the last one, list the performances of quality ladders constructed using bitrate prediction methods. The last row of the Table shows the performance of quality ladders constructed by predicting cross-over VMAF points.
From Table IX and Table X, it is evident that per-shot bitrate and quality ladders constructed using VIF feature sets yield significant gains in terms of both quality and bitrates against fixed bitrate ladders [3]. It may also be observed that the bitrate and quality ladders constructed using the ensemble of low-level features and VIF features provide even better BD-Rate and BD-VMAF gains against the fixed bitrate ladder [3], as compared to either using only low-level features or VIF features. This is also reflected in their performance against the reference bitrate ladders, as they obtain reduced BD-rate and BD-VMAF losses as compared with the non-ensemble methods. Interestingly, the performance of ladders constructed using only low-level features was better than the performance of ladders constructed using only VIF features.
All these methods learn to predict bitrate/quality ladders based on video content, which can be understood based on the performance of ladders constructed only using metadata. The metrics $\text{f}_{75}$, $\text{f}_{50}$, and $\text{f}_{25}$ provide insights into the relative performances of bitrate and quality ladders, including the reference bitrate ladder. For each method, a high value of $\text{f}_{i}$ signifies that a large proportion of test samples, encompassing both validation and test datasets, exhibit a BD-Rate and BD-VMAF gains more than $i\%$ of the BD-Rate and BD-VMAF gains of the reference bitrate ladder against the fixed bitrate ladder [3].
@@
One of the salient differences between per-shot bitrate and quality ladders constructed by predicting quality and bitrate (respectively) is that predicted per-shot bitrate ladders struggle at low bitrates and include fewer low resolutions, while the predicted per-shot quality ladders tend to struggle at higher qualities and include fewer high resolutions. This can be ascribed to errors in the ML models during prediction. For this reason, the BD-Rate gains demonstrated by the quality ladders were slightly higher than the BD-Rate gains of bitrate ladders, against a fixed bitrate ladder, since the quality ladders contain fewer high resolutions.
The per-shot bitrate and quality ladders constructed by predicting cross-over bitrates and cross-over VMAF points using low-level features $\text{LLF}_{1}$ achieved the largest gains and the least losses against the fixed bitrate ladder [3] and the reference bitrate ladder, respectively, over a substantial fraction of the test samples. Their performances were close to that of ladders constructed using true cross-over points. The bitrate ladders constructed using features $\text{LLF}_{2}$, $\text{VIFF}_{7}$ demonstrated the best performance amongst the quality prediction methods, next to bitrate ladders constructed using cross-over bitrates. Similarly, quality ladders constructed using features $\text{LLF}_{3}$, $\text{VIFF}_{9}$ demonstrated the best performance amongst the bitrate prediction methods, outperforming quality ladders constructed using cross-over VMAF points. The right selection of training and test videos, coupled with careful consideration of experimental settings, these techniques hold promise for delivering remarkable performance. Fig. 7 shows examples of convex hulls constructed using best-performing bitrate ladders, quality ladders, and fixed and reference bitrate ladders.
V-I Complexity
@@
A significant challenge associated with the utilization of the Dynamic Optimizer [6] is its substantial computational demand. Construction of a convex hull for an individual shot requires compressing the shot at many resolutions and rate control settings. Subsequently, each compressed shot must be upsampled to its original resolution to compute VMAF. This procedure imposes a considerable time and resource burden. As discussed earlier, a variety of techniques have been proposed to optimize this procedure. In our experiments, we utilized an AMD Ryzen 9 5950X and reported the total execution time, which includes both user and system time. The execution time for compression, as well as the combined execution time for upsampling and quality estimation, varied depending on the resolution and the Constant Rate Factor (CRF) used during video encoding. As a result, we present the mean and maximum observed execution times.
@@
While utilizing ffmpeg with the libx265 codec and a medium preset, we observed an approximate mean execution time of 0.42 minutes and a maximum execution time of 5.25 minutes for compression. Regarding the combined execution time of upsampling and quality estimation, we observed a mean of approximately 0.67 minutes and a maximum of around 0.94 minutes. Consequently, the construction of a convex hull via exhaustive encoding would require approximately 150 minutes on average, given our experimental settings, while direct construction of the reference bitrate ladder using bitrate steps would demand around 85 minutes on average. We also observed that the execution time for computation of the low-level feature set $\text{LLF}_{1}$ required about 6 minutes, while calculating the VIF features needed about 40 minutes. The execution time consumed by the Extra-Trees regressor was approximately 0.1 minutes per video.
@@
VI Conclusion and Future Work

### v2

@@ IV-J Constructing Per-Shot Bitrate and Quality Ladders
$q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 15.311/15.217-15.311/15.217 $3.042/3.383$ $12.546/14.353$ 2.422/2.534-2.422/2.534 $0.767$ $0.667$ $0.533$
$\text{LLF}_{3}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.546/16.722-20.546/16.722 $\textbf{4.599}/4.107$ $0.615/3.546$ 0.050/0.786-0.050/0.786 $1.000$ $0.900$ $0.900$
$\text{VIFF}_{1}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 18.337/17.154-18.337/17.154 $4.064/4.225$ $3.274/5.763$ 0.513/1.069-0.513/1.069 $0.833$ $0.767$ $0.600$
$\text{VIFF}_{2}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 17.300/17.353-17.300/17.353 $3.728/4.149$ $4.524/5.619$ 0.747/1.024-0.747/1.024 $0.833$ $0.733$ $0.533$
$\text{VIFF}_{3}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 17.905/16.652-17.905/16.652 $3.905/4.149$ $3.829/4.074$ 0.597/0.715-0.597/0.715 $0.867$ $0.800$ $0.633$
$\text{VIFF}_{4}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 18.344/16.510-18.344/16.510 $4.201/4.197$ $3.070/5.642$ 0.425/0.963-0.425/0.963 $0.900$ $0.867$ $0.667$
$\text{VIFF}_{5}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 17.584/17.344-17.584/17.344 $3.880/4.229$ $4.106/5.560$ 0.637/1.051-0.637/1.051 $0.733$ $0.700$ $0.533$
$\text{VIFF}_{6}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 18.932/16.229-18.932/16.229 $4.265/4.124$ $2.724/4.890$ 0.451/1.031-0.451/1.031 $0.933$ $0.867$ $0.700$
$\text{VIFF}_{7}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 17.426/17.025-17.426/17.025 $3.847/4.138$ $4.456/6.497$ 0.652/1.159-0.652/1.159 $0.800$ $0.800$ $0.633$
$\text{VIFF}_{8}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 17.418/16.493-17.418/16.493 $3.800/4.033$ $4.692/5.700$ 0.722/1.008-0.722/1.008 $0.833$ $0.767$ $0.533$
$\text{VIFF}_{9}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 18.452/16.350-18.452/16.350 $4.104/4.046$ $3.031/4.573$ 0.435/0.866-0.435/0.866 $0.900$ $0.800$ $0.733$
$\text{LLF}_{3}$, $\text{VIFF}_{1}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.514/16.740-20.514/16.740 $4.549/4.054$ $0.471/3.690$ 0.015/0.822-0.015/0.822 $0.933$ $0.933$ $0.867$
$\text{LLF}_{3}$, $\text{VIFF}_{2}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.138/16.605-20.138/16.605 $4.437/3.985$ $0.940/3.834$ 0.103/0.847-0.103/0.847 $1.000$ $0.900$ $0.833$
$\text{LLF}_{3}$, $\text{VIFF}_{3}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.500/16.748-20.500/16.748 $4.489/4.041$ $0.545/3.815$ 0.022/0.832-0.022/0.832 $0.967$ $0.933$ $0.833$
$\text{LLF}_{3}$, $\text{VIFF}_{4}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.727/16.557-20.727/16.557 $4.560/4.110$ $0.294/3.503$ $0.028/0.768$ $0.967$ $0.933$ $0.867$
$\text{LLF}_{3}$, $\text{VIFF}_{5}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.690/16.663-20.690/16.663 $4.568/4.005$ $0.329/3.495$ 0.000/0.793-0.000/0.793 $0.967$ $0.933$ $0.867$
$\text{LLF}_{3}$, $\text{VIFF}_{6}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.686/16.501-20.686/16.501 $4.533/3.976$ $0.438/3.477$ 0.004/0.780-0.004/0.780 $0.967$ $0.967$ $0.933$
$\text{LLF}_{3}$, $\text{VIFF}_{7}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.396/16.504-20.396/16.504 $4.506/4.122$ $0.610/3.592$ 0.030/0.789-0.030/0.789 $0.967$ $0.900$ $0.867$
$\text{LLF}_{3}$, $\text{VIFF}_{8}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.680/16.617-20.680/16.617 $4.551/4.009$ $0.310/3.503$ $0.013/0.786$ 1.000 0.967 0.900
$\text{LLF}_{3}$, $\text{VIFF}_{9}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.538/16.449-20.538/16.449 $4.530/3.972$ $0.570/3.638$ 0.030/0.818-0.030/0.818 $0.967$ $0.967$ $0.867$
$\text{LLF}_{1}$ (Cross-Over VMAFs) $\textbf{20.863}/16.728$ $4.545/4.247$ $\mathbf{-0.129}/3.304$ $\mathbf{0.102}/0.741$ $0.933$ $0.967$ $0.867$
@@
The performances of the per-shot bitrate ladders and quality ladders, constructed using the various described methods, are presented in Tables IX and X, respectively. These are compared against a fixed bitrate ladder [Fixed-Bitrate-Ladder] and reference exhaustive bitrate ladders. In Table IX, all rows, except the last one, show the performances of bitrate ladders constructed using quality prediction methods. The last row presents the performance of bitrate ladders constructed by predicting cross-over bitrates. Similarly, in Table X, all rows, barring the last one, list the performances of quality ladders constructed using bitrate prediction methods. The last row of the Table shows the performance of quality ladders constructed by predicting cross-over VMAF points. The last three columns of the Tables demonstrate the similarity of the predicted per-shot bitrate and quality ladders with the reference bitrate ladder, when comparing their performance relative to the fixed bitrate ladder. The metrics $\text{f}_{75}$, $\text{f}_{50}$, and $\text{f}_{25}$ provide insights into the relative performances of bitrate and quality ladders, including the reference bitrate ladder. For each method, a high value of $\text{f}_{i}$ signifies that a large proportion of test samples, encompassing both validation and test datasets, exhibit a BD-Rate and BD-VMAF gains more than $i\%$ of the BD-Rate and BD-VMAF gains of the reference bitrate ladder against the fixed bitrate ladder [Fixed-Bitrate-Ladder].
From Table IX and Table X, it may be observed that the per-shot bitrate and quality ladders constructed by predicting cross-over bitrates and VMAFs using the low-level features $\text{LLF}_{1}$ demonstrated excellent performance in BD-rate and BD-VMAF as compared to the fixed bitrate ladder [Fixed-Bitrate-Ladder] and the reference bitrate ladder. The per-shot bitrate ladders constructed using cross-over bitrates demonstrated a mean BD-rate loss of $1.342\%$ and a mean BD-VMAF loss of 0.124-0.124 versus the reference bitrate ladder. Similarly, the per-shot quality ladders showed a mean BD-rate loss of $0.129\%$ and a mean BD-VMAF gain of $0.102$ against the reference bitrate ladder. The negative value BD-rate against the reference bitrate ladder was likely due to noise introduced by sampling from the convex hull. Their performance is close to that of ladders constructed using true cross-over points.
From Table IX and Table X, it is evident that per-shot bitrate and quality ladders constructed using content-dependent video features extracted from the source video demonstrated improved mean BD-Rate and BD-VMAF performance relative to the fixed bitrate ladder [Fixed-Bitrate-Ladder] and the reference bitrate ladder. This is evident when the performances of our method is compared to that of ladders constructed using regressors trained on metadata. These results demonstrate the effectiveness of content-adaptive features, while also establishing a baseline against which the efficacy of video features can be determined. Regressors trained on the VIF features sets yielded significant gains in terms of both quality and bitrates, as compared to fixed bitrate ladders [Fixed-Bitrate-Ladder]. Similar to the high correlations observed between true and predicted VMAF/bitrates, in most cases regressors trained on VIF features extracted along eigen-vectors demonstrated better mean BD-Rate and BD-VMAF performance than regressors trained on features from scales and subbands. The addition of temporal features such as mean average luminance differences, and VIF features extracted on frame differences, also improved the accuracies of predicted bitrate and quality ladders. Although their performances against the reference bitrate ladder is not excellent, bitrate/quality ladders constructed using regressors trained on VIF features showed more than 50% gains of the reference bitrate ladder on approximately 70% of samples in the test dataset. The regressors trained on the $\text{VIFF}_{9}$ features demonstrated the best performance in terms of BD-metrics against the fixed bitrate ladder and the reference bitrate ladder.
Although, they achieved slightly lower correlations than did regressors trained on VIF features, the performances of per-shot ladders constructed using regressors trained on low-level features yielded better mean BD-Rate and BD-VMAF performance against the fixed bitrate ladder and the reference bitrate ladders. While the correlation gains obtained against the true VMAF/bitrates were significant, it is worth noting that the actual performances of bitrate/quality ladders is determined by the relative ordering of the predicted quality/bitrates at each resolution. Consequently, the representation of video content provided by the low-level features likely contributed to improved performance of bitrate/quality ladders constructed using these low-level features. Tables XIV and XV show the performances of per-shot bitrate and quality ladders constructed using subsets of the low-level features. The mean BD-rate losses against the reference bitrate ladders demonstrated by subsets of the low-level features was about 2% lower than the losses delivered using best performing VIF feature set ($\text{VIFF}_{9}$). Further, the performances of ladders constructed using the $\text{LLF}_{2}$ and $\text{LLF}_{3}$ features were much better than the performances of ladders constructed using subsets of the low-level features.
The regressors that were trained on the ensemble of low-level features and VIF features demonstrated better performance than either using only low-level features or VIF features. It may be observed that these methods produce mean BD-rate gains of at least 20% and BD-VMAF gains of at least 4 against the fixed bitrate ladder. Based on closeness, the per-shot bitrate ladders constructed using the ensemble of these features were only similar or slightly better than the per-shot bitrate ladders constructed using low-level features. Interestingly, the mean BD-rate and BD-VMAF values were slightly better for the per-shot bitrate ladders constructed using low-level features, on most of the feature sets. In most cases, the per-shot quality ladders constructed using the ensemble of low-level features and VIF features yielded outcomes that were closer to the reference bitrate ladder, with better mean BD-rate and BD-VMAF values, than those constructed using low-level features. These results show the effectiveness of using the ensemble of low-level features and VIF features, when constructing per-shot bitrate and quality ladders. Based on mean BD-rate and BD-VMAF values and closeness, the per-shot bitrate and quality ladders constructed using low-level features $\text{LLF}_{2}$, $\text{VIFF}_{6}$, and $\text{LLF}_{3}$, $\text{VIFF}_{8}$, respectively, delivered the best performance.
@@
One of the salient differences between per-shot bitrate and quality ladders constructed by predicting quality and bitrate (respectively) is that predicted per-shot bitrate ladders struggle at low bitrates and include fewer low resolutions, while the predicted per-shot quality ladders tend to struggle at higher qualities and include fewer high resolutions. This can be ascribed to errors in the ML models during prediction. For this reason, the BD-Rate gains demonstrated by the quality ladders were slightly higher than the BD-Rate gains of bitrate ladders, against a fixed bitrate ladder, since the quality ladders contain fewer high resolutions. The right selection of training and test videos, coupled with careful consideration of experimental settings, these techniques hold promise for delivering remarkable performance. Fig. 12 shows examples of convex hulls constructed using best-performing bitrate ladders, quality ladders, and fixed and reference bitrate ladders.
IV-K Complexity
TABLE XI: Approximate execution times of the various processes involved in the construction of the per-shot ladders using various methods.
Process Language Mean Execution Time (in minutes) Max Execution Time (in minutes)
Downsampling and Compression C 0.42 5.25
Upsampling and VMAF C 0.67 0.94
GLCM Python 2.40 -
TC Python 1.90 -
SI, TI, CF Python 1.20 -
CTI, CI Python 0.60 -
Texture-DCT Python 0.80 -
$\text{VIFF}_{3}$ Python 20.00 -
$\text{VIFF}_{9}$ Python 40.00 -
$\text{LLF}_{1}$ Python 6.00 -
Extra-Trees Regressor Python 0.10 -
@@ IV-K Complexity
A significant challenge associated with the utilization of the Dynamic Optimizer [Dynamic-Optimizer, DO] is its substantial computational demand. Construction of a convex hull for an individual shot requires compressing the shot at many resolutions and rate control settings. Subsequently, each compressed shot must be upsampled to its original resolution to compute VMAF. This procedure imposes a considerable time and resource burden. As discussed earlier, a variety of techniques have been proposed to optimize this procedure. In our experiments, we utilized an AMD Ryzen 9 5950X and reported the total execution time, which includes both user and system time. The execution time for compression, as well as the combined execution time for upsampling and quality estimation, varied depending on the resolution and the Constant Rate Factor (CRF) used during video encoding. As a result, we present the mean and maximum observed execution times.
@@
Table XI shows the mean and maximum execution times of various procedures involved in the construction of per-shot bitrate/quality ladders using our proposed method, versus constructing reference bitrate ladders using exhaustive encoding. When utilizing ffmpeg with the libx265 codec and a medium preset, we observed an approximate mean execution time of 0.42 minutes and a maximum execution time of 5.25 minutes for compression. We observed a mean combined execution time of approximately 0.67 minutes, with a maximum of around 0.94 minutes, for the upsampling and quality estimation processes. Consequently, the construction of a reference bitrate ladder sampled from a convex hull via exhaustive encoding would require approximately 150 minutes on average, given our experimental settings. We also observed that the execution time to compute the low-level feature set $\text{LLF}_{1}$ required about 6 minutes, while calculating the VIF features $\text{VIFF}_{9}$ needed about 40 minutes. The execution time consumed by the Extra-Trees [Extra-Trees] regressor was approximately 0.1 minutes per video. Table XI also shows the execution times to compute various subsets of low-level features and VIF features.
@@
TABLE XII: Pearson correlation coefficients between true VMAF and VMAF predicted using various subsets of low-level features at each resolution of the validation and test datasets.
Features 2160p 1440p 1080p 720p 540p 432p
GLCM, TC 0.529 0.595 0.631 0.646 0.627 0.588
SI, TI 0.540 0.606 0.634 0.646 0.618 0.593
CTI, CF, CI 0.440 0.494 0.527 0.551 0.500 0.385
DCT-Texture, Bitrate-DCT-Texture 0.619 0.673 0.690 0.674 0.616 0.522
$\text{LLF}_{2}$ 0.580 0.642 0.668 0.677 0.643 0.602
TABLE XIII: Pearson correlation coefficients between true bitrate and bitrate predicted using various subsets of low-level features at each resolution of the validation and test datasets.
Features 2160p 1440p 1080p 720p 540p 432p
GLCM, TC 0.538 0.569 0.581 0.582 0.554 0.517
SI, TI 0.615 0.647 0.663 0.678 0.671 0.655
CTI, CF, CI 0.500 0.525 0.536 0.520 0.463 0.394
DCT-Texture, VMAF-DCT-Texture 0.601 0.652 0.684 0.703 0.701 0.688
$\text{LLF}_{3}$ 0.565 0.598 0.617 0.626 0.613 0.595
TABLE XIV: Means and standard deviations of BD-metrics, and closeness of each model’s predicted per-shot bitrate ladders against fixed and reference bitrate ladders on the validation and test datasets, trained on various subsets of low-level features.
BD-Rate (in %) BD-VMAF BD-Rate (in %) BD-VMAF
GLCM, TC, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 19.860/17.132-19.860/17.132 $4.116/3.887$ $1.509/4.051$ 0.203/0.705-0.203/0.705 $0.933$ $0.833$ $0.767$
SI, TI, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 19.044/16.788-19.044/16.788 $4.136/4.090$ $2.046/5.831$ 0.248/0.987-0.248/0.987 $0.900$ $0.800$ $0.733$
CTI, CF, CI, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 17.498/21.019-17.498/21.019 $3.736/4.789$ $2.824/6.079$ 0.596/1.514-0.596/1.514 $0.867$ $0.800$ $0.733$
DCT-Texture, Bitrate-DCT-Texture, $\log_{2}(b)$, $\frac{w}{3840}$ 19.464/18.260-19.464/18.260 $4.066/4.210$ $1.197/4.246$ 0.191/0.815-0.191/0.815 $0.867$ $0.800$ $0.767$
$\text{LLF}_{2}$, $\log_{2}(b)$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.429/16.747-20.429/16.747 $4.323/4.092$ $0.439/4.791$ 0.016/1.004-0.016/1.004 $0.967$ $0.933$ $0.767$
TABLE XV: Means and standard deviations of BD-metrics, and closeness of each model’s predicted per-shot quality ladders against fixed and reference bitrate ladders on the validation and test datasets, trained on various subsets of low-level features.
BD-Rate (in %) BD-VMAF BD-Rate (in %) BD-VMAF
GLCM, TC, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.204/16.671-20.204/16.671 $4.462/4.117$ $0.975/3.653$ 0.118/0.776-0.118/0.776 $0.967$ $0.933$ $0.833$
SI, TI, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 19.595/16.347-19.595/16.347 $4.300/4.062$ $1.487/3.430$ 0.191/0.741-0.191/0.741 $0.967$ $0.900$ $0.800$
CTI, CF, CI, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 18.816/16.868-18.816/16.868 $4.105/4.156$ $2.545/5.125$ 0.368/1.095-0.368/1.095 $0.867$ $0.800$ $0.700$
DCT-Texture, VMAF-DCT-Texture, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 19.904/16.712-19.904/16.712 $4.454/4.072$ $1.184/4.086$ 0.152/0.877-0.152/0.877 $0.933$ $0.833$ $0.767$
$\text{LLF}_{3}$, $q/100$, $\frac{w}{3840}$, $\frac{h}{3840}$ 20.546/16.722-20.546/16.722 $4.599/4.107$ $0.615/3.546$ 0.05/0.786-0.05/0.786 $1.000$ $0.900$ $0.900$
V Conclusion and Future Work